# SMART: Self-Generating and Self-Validating Multi-Dimensional Assessment for LLMs’ Mathematical Problem Solving

### v1

@@ Abstract
Large Language Models have achieved remarkable results on a variety of mathematical benchmarks. However, concerns remain as to whether these successes reflect genuine mathematical reasoning or superficial pattern recognition. Common evaluation metrics, such as final answer accuracy, fail to disentangle the underlying competencies involved, offering limited diagnostic value. To address these limitations, we introduce SMART: a Self-Generating and Self-Validating Multi-Dimensional Assessment Framework. SMART decomposes mathematical problem solving into four distinct dimensions: understanding, reasoning, arithmetic, and reflection & refinement. Each dimension is evaluated independently through tailored tasks, enabling interpretable and fine-grained analysis of LLM behavior. Crucially, SMART integrates an automated self-generating and self-validating mechanism to produce and verify benchmark data, ensuring both scalability and reliability. We apply SMART to 21 state-of-the-art open- and closed-source LLMs, uncovering significant discrepancies in their abilities across different dimensions. Our findings demonstrate the inadequacy of final answer accuracy as a sole metric and motivate a new holistic metric to better capture true problem-solving capabilities. Code and benchmarks will be released upon acceptance.
SMART: Self-Generating and Self-Validating Multi-Dimensional Assessment for LLMs’ Mathematical Problem Solving
Yujie Hou, Ting Zhang, Mei Wang, Xuetao Ma, Hua Huangthanks:  Corresponding author: Hua Huang School of Artificial Intelligence, Beijing Normal University, Beijing, China {houyujie, maxuetao}@mail.bnu.edu.cn, {tingzhang,wangmei1, huahuang}@bnu.edu.cn
1 Introduction
@@
Large language models (LLMs)  Achiam et al. (2023); Wei et al. (2022) have demonstrated impressive performance across various natural language processing tasks Singhal et al. (2023); Wang et al. (2024); Jiang et al. (2024); Ma et al. (2025) and are being increasingly integrated into real-world applications, such as education, scientific computing, and decision support OpenAI (2024); Guo et al. (2025). As their adoption accelerates, it is critical to rigorously evaluate their true capabilities.
Figure 1: Following Polya’s four-step framework for human mathematical problem solving—understanding the problem, devising a plan, carrying out the plan, and looking back—we define four corresponding evaluation dimensions and design specific tasks to systematically assess the capabilities of large language models.
While numerous benchmarks have been proposed to evaluate the mathematical reasoning capabilities of LLMs  Cobbe et al. (2021); Ling et al. (2017); Patel et al. (2021); Miao et al. (2020); Koncel-Kedziorski et al. (2016), many suffer from a significant limitation: they provide insufficient interpretability regarding actual mathematical proficiency. The predominant focus on final answer accuracy is particularly problematic, as this metric reflects an aggregate outcome influenced by multiple, intertwined cognitive processes. As a result, existing benchmarks fail to effectively evaluate the distinct components of the mathematical problem-solving process Cobbe et al. (2021); Ling et al. (2017); Patel et al. (2021); Miao et al. (2020); Koncel-Kedziorski et al. (2016).
@@
Another critical deficiency stems from the static nature of most existing benchmarks. Such static evaluation sets are highly susceptible to data contamination, potentially leading to inflated performance scores and skewed assessments of LLM capabilities Oren et al. (2023); Zhu et al. (2023, 2024); Li et al. (2024a). Furthermore, as LLMs rapidly advance, they can quickly reach performance saturation on these fixed datasets. Consequently, static benchmarks may no longer accurately reflect the true capabilities of more advanced models Fodor (2025); AI@Meta (2024).
Furthermore, constructing high-quality benchmark datasets is resource-intensive, requiring significant human labor and time for annotation Mirzadeh et al. (2024); Kurtic et al. (2024). When leveraging advanced LLMs like GPT-4o for data generationLi et al. (2024b); Zhu et al. (2024), ensuring the accuracy and consistency of the generated content remains a major challenge, as LLMs may produce errors and inconsistencies. Although human annotators are often employed to review and refine LLM-generated questions and answers Zhu et al. (2024); Zheng et al. (2023), this manual intervention is costly and inherently limits scalability.
@@
Figure 2: An overview of the SMART framework for evaluating the mathematical problem-solving process. The SMART contains four distinct dimension. Each dimension is evaluated using dimension-specific task and metrics, ensuring a comprehensive assessment of the model’s problem-solving capabilities.
@@
In light of these issues, we propose a novel evaluation framework: the Self-Generating and Self-Validating Multi-Dimensional Assessment Framework (SMART). Inspired by Polya’s problem-solving theory Polya (2014), SMART deconstructs the mathematical problem-solving process into four key evaluation dimensions: understanding, reasoning, arithmetic, and reflection & refinement. Each dimension is designed to evaluate a distinct facet of problem-solving progression (Fig. 1).
To implement this multi-dimensional assessment, SMART generates dimension-specific tasks with clearly defined outputs from seed problems. This structured decomposition allows SMART to enable the model to probe distinct cognitive components of mathematical problem solving. By isolating each dimension, SMART facilitates comprehensive evaluation and yields insights into the strengths and weaknesses of LLMs, informing targeted algorithmic enhancements. The resulting evaluation dataset, SMART-Bench, is built through an integrated self-generating and self-validating mechanism that not only programmatically produces dimension-specific data but also iteratively verifies its quality. This approach ensures both scalability and reliability in benchmark construction.
We evaluate 21 recently released open- and closed-source LLMs on SMART-Bench. Our experimental results reveal that models exhibit diverse performance patterns, demonstrating varied strengths and weaknesses across different dimensions. Furthermore, our investigation pinpoints key factors influencing performance across these distinct dimensions. Finally, through an in-depth analysis of the relationship between final answer accuracy and the scores of our decomposed dimensions, we propose a new metric designed to more accurately measure the true mathematical capabilities of LLMs. Our main contributions are as follows:
@@ 1. $\bullet$
    We propose SMART, a self-generating and self-validating multi-dimensional assessment framework designed to break down the mathematical problem-solving process of LLMs into four key dimensions. This approach enables deeper insight into underlying cognitive processes beyond final answer accuracy.
@@ 2. $\bullet$
    We introduce an integrated self-generating and self-validating mechanism that produces SMART-bench dataset. This significantly reduces reliance on human annotation while ensuring the benchmark’s scalability and reliability.
@@ 2 Related Work
Mathematic benchmark. Numerous mathematical benchmarks with varying levels of difficulty have been developed to explore the upper bound of LLMs’ mathematical capabilities. These benchmarks range from grade-school-level datasets Cobbe et al. (2021), to high-school-level datasets Hendrycks et al. (2021), and extend to expert-level datasets Glazer et al. (2024). Their scope covers a broad range of mathematical domains, including geometry, number theory, and real analysis. However, despite the increasing difficulty of these benchmarks, it remains unclear whether LLMs truly comprehend mathematical concepts or merely rely on pattern-matching to generate correct answers Mirzadeh et al. (2024). Furthermore, the widespread use of these benchmarks increase the risk of data contamination, potentially inflating performance evaluations Li et al. (2024a).
Dynamic evaluation. To address these concerns, recent studies have explored dynamic evaluation approaches Zhu et al. (2023, 2024), which generate adaptive test data based on pre-defined rewriting rules. For example, GSM-Plus Li et al. (2024b) introduces eight perturbation strategies, while GSM-Symbolic Mirzadeh et al. (2024) generates diverse problem variants using symbolic templates. These approaches aim to mitigate data leakage and improve robustness in assessments. Despite these advancements, current evaluation methodologies remain limited. They predominantly focus on final-answer accuracy, which fails to capture an LLM’s reasoning process in a systematic manner. Additionally, they do not explicitly assess the diverse cognitive skills required for mathematical problem-solving, nor do they provide fine-grained insights into the underlying logical reasoning steps.
@@
These limitations underscore the need for a comprehensive, interpretable, and scalable evaluation framework that can decompose the problem-solving process, assess reasoning at multiple levels, and minimize reliance on human verification. To bridge this gap, we introduce SMART, a self-validating multi-dimensional assessment framework designed to systematically evaluate LLMs’ mathematical reasoning abilities while enhancing benchmark reliability and scalability.
3 Methodology
3.1 SMART Framework
@@
Understanding. The initial step in mathematical problem-solving is to comprehend the problem. Accordingly, the first dimension in the SMART framework assesses the ability of LLMs to clearly and comprehensively understand a given problem before attempting a solution. To rigorously evaluate this dimension, we introduce a context extraction task. This task requires the LLM to meticulously identify and categorize various informational components embedded within the problem. As illustrated in Fig. 2 (Dimension 1, Task 1), the LLM is prompted to extract and structure key information into a predefined template. This template includes categories such as: problem scenario, goal, known and unknown quantities, relationships and constraints, and irrelevant information.
@@
This structured context extraction task serves a dual purpose. At a surface level, it evaluates the LLM’s ability to summarize and extract salient information. More profoundly, it probes the depth of the LLM’s comprehension by requiring it to discern the roles and interconnections of different problem elements, thereby offering a nuanced measure of its problem understanding capabilities. This moves beyond simple keyword spotting to assess a more holistic grasp of the problem’s structure.
@@
Reasoning. Subsequent to the comprehension phase, Polya’s problem-solving heuristic emphasizes devising a coherent plan. Aligning with this, the second dimension evaluates the reasoning ability of LLMs. This is achieved through a symbolic formalization task, wherein LLMs are required to solve mathematical problems via formal symbolic expressions. This approach compels the LLM to concentrate on the inherent logical structure of the problem and the intricate relationships between its constituent elements, rather than relying on direct numerical computation.
@@
Beyond merely translating problem statements into symbolic form, this task demands the identification of solution-relevant equations and the construction of a logically consistent path to the answer. As illustrated in Fig. 2 (Dimension 2, Task 2), a highlighted symbolic assertion, absent from the original question, captures a key inferred relation, indicating the model’s ability to derive implicit knowledge. This dimension thus probes deeper reasoning by emphasizing inference and the construction of a complete logical model.
We adopt a standardized and widely used format for representing satisfiability modulo theories (SMT) problems Barrett et al. (2010). The SMT-LIB expressions produced by the models are subsequently verified using established SMT solvers such as Z3 De Moura and Bjørner (2008), which can automatically determine the satisfiability of logical formulas or compute concrete solutions. A reasoning plan expressed in SMT-LIB is considered correct if the solution derived by the solver consistently align with the ground truth of the original mathematical problem.
@@
Arithmetic. The third dimension, Arithmetic, is dedicated to assessing an LLM’s proficiency in performing accurate numerical computations. This evaluation is conducted through a numeric calculation task that presents problems exclusively in terms of numerical values and variables, requiring the execution of fundamental arithmetic operations. For instance, as illustrated in Fig.  2 (Dimension 3, Task 3), the arithmetic questions are designed to contain only basic calculation operations and clearly defined variables.
This task design isolates arithmetic skills by minimizing confounding cognitive demands such as language comprehension or complex reasoning, thereby enabling a focused and accurate evaluation of the model’s computational proficiency.
Reflection & Refinement. The final stage in Polya’s problem-solving model, "looking back," involves reviewing the solution and improving its correctness. Recent literature has demonstrated that LLMs possess a promising capacity for self-feedback, and iterative refinement Madaan et al. (2023); Shinn et al. (2023); Guo et al. (2025); Huang et al. (2022). These studies suggest that prompting LLMs to scrutinize their own outputs or follow structured verification steps lead to improved accuracy and more robust reasoning. So, it is crucial to evaluate the self-critique capability of LLMs.
Accordingly, the fourth dimension of our SMART framework, Reflection & Refinement, aims to evaluate the LLM’s ability to critically review its problem-solving process and refine its answers. This is operationalized through an error correction task, an aspect largely overlooked in prior benchmarks. As depicted in Fig. 2 (Dimension 4, Task 4), LLMs are presented with a Chain-of-Thought (CoT) solution that contains deliberately introduced errors. We introduce six distinct categories of errors into the CoT, encompassing arithmetic number error, skipped step error, hallucinatory injection error, logical order error, redundant output error, and incorrect operator error (detailed in Fig. 18). The LLM is then tasked with first identifying any flaws within the provided CoT, then correcting these identified errors, and finally generating a refined answer based on the corrected reasoning. A critical design choice is that if an LLM fails to detect any errors in the provided CoT, it does not proceed to the subsequent refinement tasks for that particular problem instance.
3.2 The Self-Generating and Self-Validating Mechanism in SMART
A core innovation of SMART lies in its integrated self-generating and self-validating mechanism, which ensures both scalable benchmark construction and the reliability of SMART-Bench data.
3.2.1 Dataset Generation
@@
To construct SMART-Bench, we begin by curating a diverse set of 2,000 seed questions, selected from seven widely used mathematical problem datasets: GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), ASDiv Miao et al. (2020), AQuA Ling et al. (2017), MAWPS Koncel-Kedziorski et al. (2016), MATH Hendrycks et al. (2021), and the AIME 2024 competition Huggingface (2024). For each seed question, we automatically generate four dimension-specific variants, each aligned with a core dimension. The resulting dataset comprises 10,000 instances, 2,000 original questions and 8,000 generated variants. Details about generation procedures are provided in Appendix B.
3.2.2 Dataset Verification
For the understanding dimension, given the strong capability of LLMs in information extraction, we rely on a robust LLM to ensure correctness. In contrast, for reasoning and arithmetic questions, manual verification by human annotators is resource-intensive and limits scalability. To overcome this, SMART incorporates a self-validating mechanism based on a neuro-symbolic framework that automatically and precisely verifies generated question variants and their answers for dimensions suitable for symbolic representation. For example, to verify the correctness of a generated SMT-LIB expression for a seed question, we employ an SMT solver to compute the symbolic formula’s answer automatically. We then compare the solver’s output with the known seed answer, iterating this process until the generated SMT-LIB expression aligns with the expected result. Further details are provided in Appendix C. For the Reflection & Refinement dimension, ground truth is guaranteed since errors are intentionally produced.
To further assess the robustness of LLM capabilities, SMART employs a suite of rewriting rules to create augmented versions of the dimension-specific tasks. The specifics of these rewriting strategies and the results of associated experiments are detailed in Sec. 4.3 and 4.4. The self-validation methodology remains effective for ascertaining the correctness of these newly formulated, augmented questions, even when their answers differ from those of the original problems. This capability is crucial for maintaining the integrity of dynamically generated test instances. A comprehensive description of the procedures for the self-validation of augmented data is available in Appendix C.
4 Experiments
4.1 Experimental Setup
We evaluate 21 recent open-source and closed-source LLMs using our SMART evaluation framework. The open-source models include Phi4 Abdin et al. (2024), Gemma3 Team et al. (2025), GLM4 GLM et al. (2024), Qwen2.5 Team (2024), Qwen3 Yang et al. (2025), Llama3 AI@Meta (2024), Llama4 Meta (2025), and Mistral MistralAITeam (2024). The closed-source models assessed are GPT-4o Achiam et al. (2023), GPT-4.1 OpenAI (2025a), o4-mini, o3 OpenAI (2025b), DeepSeek-V3 Liu et al. (2024), DeepSeek-R1 Guo et al. (2025), Claude3.5 anthropic (2024), Claude3.7 anthropic (2025), and Gemini2.5 Google (2024). The temperature is set to 0 to ensure deterministic outputs. We employ an LLM-as-a-judge approach Zheng et al. (2023) to evaluate the understanding dimension. Tasks for the other dimensions primarily utilize accuracy-based metrics. For all dimension-specific tasks, we adopt a three-shot prompting strategy. Detailed descriptions of the evaluation metrics and prompt designs for each of the four SMART dimensions are provided in Appendix D.2.
@@ 4.1 Experimental Setup
Model Final Answer Understanding Reasoning Arithmetic Reflection&Refinement Reflection Refinement
ACC@Fi LLM@Un ACC@Re ACC@Ar ACC@RR ACC@R-t ACC@R-m
@@
Phi4-14B 89.55 93.17 64.12 96.45 22.06 26.75 82.34
Gemma3-27B 54.62 94.47 40.81 39.94 24.23 32.55 74.35
GLM4-32B 54.13 94.39 55.92 36.62 22.85 28.50 88.57
Qwen2.5-72B 41.05 95.21 62.53 35.55 33.85 35.75 94.69
Qwen3-32B 93.75 93.19 86.14 98.50 54.75 56.70 96.52
Qwen3-30B-A3B 92.85 93.29 85.20 98.15 43.89 58.25 75.36
Llama3.1-8B 45.15 88.92 9.05 62.15 3.30 8.80 37.46
Llama3.3-70B 88.95 95.04 49.26 94.72 40.30 41.75 96.53
Llama4-Scout-17B-16E 91.24 94.48 62.65 97.11 29.04 49.5 58.67
Mistral-Small-3.1-24B 77.55 94.48 28.96 49.55 24.55 36.65 66.98
Mistral-Large-123B 62.25 94.75 48.95 43.74 43.12 50.61 81.62
@@
GPT-4o 57.65 95.32 64.45 45.64 42.60 45.30 94.04
GPT-4.1 59.31 95.41 73.29 45.71 59.90 63.80 93.89
o4-mini 93.75 95.22 90.72 98.25 55.05 64.61 85.22
o3 91.44 95.17 89.12 98.45 71.51 78.62 90.97
DeepSeek-V3 93.43 94.92 65.25 98.35 68.65 71.10 96.55
DeepSeek-R1 97.92 94.88 93.61 98.15 62.50 69.80 89.54
Claude3.5-Sonnet 92.61 94.69 80.85 97.61 33.05 65.85 50.19
Claude3.7-Sonnet 93.53 95.09 74.18 98.11 57.43 66.35 86.51
Gemini2.5-Flash-Preview 93.33 94.25 89.72 98.12 35.85 50.35 71.20
Gemini2.5-Pro-Preview 94.15 94.23 91.45 97.95 54.80 76.90 71.26

### v3

@@ Abstract
Large Language Models (LLMs) have achieved remarkable results on a variety of mathematical benchmarks. However, concerns remain as to whether these successes reflect genuine reasoning or superficial pattern recognition. Common evaluation methods, which focus on the either the final answer or the reasoning process, fail to assess the entire problem-solving procedure. To address these limitations, we introduce SMART: a Self-Generating and Self-Validating Multi-Dimensional Assessment Framework, together with its corresponding benchmark, SMART-Bench. SMART decomposes the entire problem solving process into four distinct cognitive dimensions: Understanding, Reasoning, Arithmetic, and Reflection & Refinement. Each dimension is evaluated independently through tailored tasks, enabling interpretable and fine-grained analysis of LLM behavior. We apply SMART to 21 state-of-the-art open- and closed-source LLMs, uncovering significant discrepancies in their abilities across different dimensions. Our findings reveal genuine weaknesses in current LLMs and motivate a new metric, the All-Pass Score, to better capture true problem-solving capabilities. Code and benchmarks will be released upon acceptance.
Introduction
@@
Large language models (LLMs)  (Achiam et al. 2023; Wei et al. 2022) have demonstrated impressive performance across various natural language processing tasks (Singhal et al. 2023; Wang et al. 2024; Jiang et al. 2024; Ma, Jiang, and Huang 2025) and are being increasingly integrated into real-world applications, such as education, scientific computing, and decision support (OpenAI 2024; Guo et al. 2025). With the increasing adoption of LLMs, it is critical to rigorously evaluate their true capabilities.
Figure 1: Comparison between SMART and existing methods. Final answer-based evaluations assess only the correctness of the final outcome. Reasoning process-based evaluations focus on identifying errors within the solution path. Our proposed SMART method decomposes the entire problem-solving process into four distinct dimensions, each evaluated through targeted tasks. This enables a more fine-grained and systematic assessment of LLMs’ mathematical problem-solving abilities.
Mathematical questions are widely used as proxy tasks to evaluate the cognitive abilities of LLMs. Although numerous mathematical benchmarks have been proposed, several important limitations persist. A primary issue is that current datasets do not holistically assess the entire problem-solving process, which involves multiple, intertwined cognitive dimensions. For example, widely used benchmarks such as GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021) focus primarily on final answer accuracy, while others—such as MR-GSM8K (Zeng et al. 2024a)—target specific aspects like the reasoning process. A comprehensive process-level evaluation is essential for measuring the true capabilities of LLMs in mathematical problem solving.
@@
Another critical deficiency stems from the static nature of most existing benchmarks. Such static evaluation sets are highly susceptible to data contamination, potentially leading to inflated performance scores and skewed assessments of LLM capabilities (Oren et al. 2023; Zhu et al. 2023, 2024; Li et al. 2024a). Furthermore, as LLMs rapidly advance, they can quickly reach performance saturation on these fixed datasets. As a result, static benchmarks may no longer accurately reflect the true capabilities of more advanced models (Fodor 2025; AI@Meta 2024).
Furthermore, constructing high-quality mathematical benchmarks demands considerable human effort and annotation resources. For instance, miniF2F (Zheng, Han, and Polu 2021) and FIMO (Liu et al. 2023) manually formalize IMO-level problems with carefully curated formal statements or solutions. Similarly, MR-BEN (Zeng et al. 2024b) involves costly step-level annotations. As the number of samples increases, the annotation cost rises dramatically, significantly limiting the scalability of such benchmarks.
@@
Figure 2: An overview of the SMART framework for evaluating the mathematical problem-solving process. The SMART contains four distinct dimension. Each dimension is evaluated using dimension-specific task and metrics, ensuring a comprehensive assessment of the model’s problem-solving capabilities. The yellow-highlighted symbolic assertion illustrates an inferred condition that is not explicitly stated in the original problem.
@@
In light of these limitations, we propose a novel evaluation framework: the Self-Generating and Self-Validating Multi-Dimensional Assessment Framework (SMART), along with a new metric called the All-Pass Score. Inspired by Polya’s problem-solving theory (Polya 2014), SMART decomposes the mathematical problem-solving process into four key evaluation dimensions: Understanding, Reasoning, Arithmetic, and Reflection & Refinement (Fig. 1). This evaluation paradigm enables a comprehensive assessment of the entire problem-solving process—from initial problem comprehension, to planning reasoning steps, performing numerical computations, and finally checking the solution.
To implement this multi-dimensional assessment, we introduce a self-generating and self-validating mechanism that dynamically constructs dimension-specific tasks from seed problems, forming our benchmark SMART-Bench. These dimension-specific tasks are novel to LLMs and thus help alleviate data contamination. We adopt a neuro-symbolic framework (Barrett et al. 2010; De Moura and Bjørner 2008) to iteratively verify the correctness of the generated questions, significantly reducing the need for manual annotation. Thus, our approach ensures the scalability and reliability of benchmark construction.
We evaluate 21 recently released open- and closed-source LLMs on SMART-Bench. Experimental results demonstrate that even the most advanced models perform poorly under the proposed All-Pass Score, underscoring the challenging nature of our benchmark. Notably, the Reasoning and Reflection dimensions emerge as the primary bottlenecks in mathematical problem solving. Additionally, SMART serves as a diagnostic tool, identifying which dimensions are responsible for performance drops under dynamic evaluation. Furthermore, we find that targeted improvements in specific weak dimensions can lead to substantial overall gains in mathematical capability—for example, increasing the final answer accuracy of Qwen2.5-72B by 11.77%. Our main contributions are as follows:
@@ 1. $\bullet$
    We propose the SMART evaluation paradigm along with a novel All-Pass Score metric to enable fine-grained and comprehensive assessment of a model’s mathematical problem-solving abilities.
@@ 2. $\bullet$
    We introduce an integrated self-generating and self-validating mechanism for constructing SMART-Bench. The self-generating component creates novel tasks from seed problems to mitigate data contamination, while the self-validating component verifies correctness to reduce human annotation. Together, they ensure the scalability and reliability of the benchmark.
@@ 3. $\bullet$
Mathematic benchmark Numerous mathematical benchmarks with varying levels of difficulty have been developed to explore the upper bound of LLMs’ mathematical capabilities. These benchmarks range from grade-school-level datasets (Cobbe et al. 2021), to high-school-level datasets (Hendrycks et al. 2021), and extend to expert-level datasets (Glazer et al. 2024). Their scope covers a broad range of mathematical domains, including geometry, number theory, and real analysis. However, despite their increasing difficulty, these benchmarks primarily adopt an outcome-based evaluation approach, making it unclear whether LLMs genuinely understand mathematical concepts or simply rely on pattern-matching to produce correct answers (Mirzadeh et al. 2024). To address this, ProcessBench (Zheng et al. 2025) and PRMBench (Song et al. 2025) have been proposed to enable process-level evaluation by identifying erroneous steps in the model’s mathematical reasoning. Nevertheless, these process-level benchmarks still fall short of evaluating the entire spectrum of human mathematical thinking, as they do not comprehensively assess all stages of the problem-solving process
Dynamic evaluation The widespread use of benchmarks increase the risk of data contamination, potentially inflating performance evaluations (Li et al. 2024a). To address these concerns, recent studies have explored dynamic evaluation approaches (Zhu et al. 2023, 2024), which generate adaptive test data based on pre-defined rewriting rules. GSM-Plus (Li et al. 2024b) and GSM-Symbolic (Mirzadeh et al. 2024) generates problem variants from seed questions. These approaches aim to mitigate data leakage and improve robustness in evaluations. However, manually annotating the newly generated questions remains highly labor-intensive and costly. This motivates the need for automated mechanisms that can generate and verify evaluation data at scale.
@@
These limitations underscore the need for a framework that can comprehensively assess the entire problem-solving process, provide fine-grained and interpretable analyses, and reduce the cost of benchmark construction. To address this gap, the SMART framework is designed to systematically evaluate the mathematical reasoning abilities of LLMs.
Dimension Input question Input source Output Ground truth Ground truth source Task evaluator
Understanding Seed question / Structured information Structured information
GPT-4.1
+ human verification
LLM-as-a-judge
Reasoning Seed question / SMT-LIB
Final answer of
seed question
Z3 Solver and
Rule-match
Arithmetic
Arithmetic
notation question
SMT-LIB+GPT-4.1
+ human verification
Answer of
arithmetic question
Final answer of
seed question
/ Rule-match
Reflection
CoT solution with
injected errors
Rule Error types Error types / Rule-match
Refinement
CoT solution with
injected errors
Rule Corrected CoT
Final answer of
seed question
/ Rule-match
Table 1: Overview of Dimension-specific Task Design and Ground truth in SMART
Methodology
Evaluation Dimensions
Understanding
@@
Understanding the problem is the first step in mathematical problem solving. Accordingly, the first dimension of SMART evaluates an LLM’s understanding ability through an information extraction task, which requires the model to accurately identify and categorize key components within the problem statement. As illustrated in Fig. 2 Dimension 1, the LLM is prompted to extract and organize this information into a predefined template. The template includes categories such as the problem scenario, goal, known and unknown quantities, relationships and constraints, and irrelevant information.
@@
This structured context extraction task serves a dual purpose. At a surface level, it evaluates the LLM’s ability to summarize and extract salient information. More profoundly, it probes the depth of the LLM’s comprehension by requiring it to discern the roles and interconnections of different elements, thereby offering a nuanced measure of its problem understanding capabilities.
Reasoning
@@
Following the comprehension phase, the next stage involves devising a reasoning plan. Thus, the second dimension of SMART evaluates the reasoning ability of LLMs through a symbolic formalization task, in which models are required to solve mathematical problems using formal symbolic expressions in SMT-LIB format (Barrett et al. 2010). This approach compels the model to focus on the underlying logical structure of the problem and the intricate relationships among its components, rather than relying on direct numerical computation. Although the SMT-LIB format is seldom used in LLMs, experiments presented in the Appendix demonstrate that the primary challenge of the Reasoning task lies in generating a correct reasoning path, rather than translating natural language into formal language.
@@
The reasoning task extends beyond merely translating problem statements into symbolic form; it requires LLMs to identify solution-relevant equations and construct a logically coherent path to the answer. As illustrated in Fig. 2 (Dimension 2), the highlighted symbolic assertion is not explicitly provided as a known condition in the original question. Instead, LLMs must infer this key constraint to solve the problem, thereby demonstrating their ability to derive implicit knowledge.
Arithmetic
@@
The third dimension, Arithmetic, is dedicated to assessing an LLM’s proficiency in performing accurate numerical computations. This evaluation is conducted through a numeric calculation task that presents problems exclusively in terms of numerical values and variables, requiring the execution of fundamental arithmetic operations, as illustrated in Fig.  2 Dimension 3. This experimental setup isolates arithmetic skills from other cognitive demands—such as language comprehension or complex reasoning—thereby enabling a focused and precise evaluation of the model’s arithmetic capabilities.
Reflection & Refinement The fourth dimension of SMART is Reflection & Refinement, which assesses the LLM’s capacity for self-critique and iterative improvement. As illustrated in Fig. 2, Dimension 4, the model is presented with a question and its chain-of-thought (CoT) solution, and is tasked with identifying potential errors in CoT. It then revise the errors and produce an corrected CoT. A key design feature is that if the LLM fails to detect all errors in the initial solution, it is not allowed to proceed to the refinement stage for that instance.
Data Curation
We adopt a self-generating approach to construct SMART-Bench from a set of seed questions. Tab. 1 provides a summary of SMART-Bench task, ground truth, and evaluation for each dimension in SMART.
Seed Question Collection
@@
We begin by collecting a diverse set of 2,000 seed questions from seven widely-used mathematical problem datasets: GSM8K (Cobbe et al. 2021), SVAMP (Patel, Bhattamishra, and Goyal 2021), ASDiv (Miao, Liang, and Su 2020), AQuA (Ling et al. 2017), MAWPS (Koncel-Kedziorski et al. 2016), MATH (Hendrycks et al. 2021), and the AIME 2024 (Huggingface 2024). These seed questions serve as the foundation for generating dimension-specific evaluation tasks.
Test Case Generation
Each dimension in SMART-Bench is designed with a distinct test case tailored to evaluate specific aspects of LLMs. For the Understanding and Reasoning dimensions, the input questions are identical to the original seed problems, but each targets to generate different outputs as shown in Tab.1. For the Understanding task, we use structured information extracted by GPT-4.1 and further verified by human annotators as the ground truth.
Directly converting a seed question into an arithmetic notation problem is challenging, as it requires simplifying complex natural language into structured mathematical operations while preserving logical relationships between variables. Therefore, for the Arithmetic dimension, we adopt a two-stage process: seed questions are first formalized into SMT-LIB representations using GPT-4.1 to capture the core logic, and these formal expressions are then translated into arithmetic problems, also with GPT-4.1.
For the Reflection dimension, the input consists of the original question paired with a chain-of-thought (CoT) solution containing deliberately injected errors. We define six categories of errors: arithmetic inaccuracies, omitted steps, hallucinated content, logical disorder, redundancy, and operator misuse. All errors are generated based on predefined rules to simulate flawed reasoning processes. Detailed descriptions of these CoT errors are provided in the Appendix.
The Refinement task takes the same input as Reflection task but ask to generate the corrected CoT.
Quality Control
As shown in Tab. 1, only the input questions for the Arithmetic task and the ground truth for the Understanding task are generated by LLMs and thus require additional verification. To ensure the quality of SMART-Bench, we implement a self-validating mechanism comprising both neuro-symbolic and human verification procedures. This mechanism automatically identifies and filters out low-quality samples. If a sample fails the validation process, it is iteratively regenerated until it satisfies the required quality standards.
Neuro-symbolic verification
We employ neuro-symbolic verification to ensure the correctness of the SMT-LIB expressions used in the Arithmetic dimension. Directly comparing generated SMT-LIB with ground-truth expressions is challenging. Instead, we leverage the Z3 solver (De Moura and Bjørner 2008) to automatically compute the result of a symbolic formula and compare it against the ground-truth answer from the original seed question, as a correct SMT-LIB expression should yield the same answer as its seed question. This generation–validation process is repeated until the SMT-LIB expression produces the correct answer.
@@
Model Understanding Reasoning Arithmetic $R$ Reflection Refinement Final Answer All-Pass Score
LLM@Un ACC@Re ACC@Ar ACC@$R$ ACC@R-t ACC@R-m ACC@Fi ACC@All
@@
Phi4-14B 93.41 64.12 96.45 22.06 26.75 82.34 89.55 20.77
Gemma3-27B 94.31 40.81 39.94 24.23 32.55 74.35 54.62 11.96
GLM4-32B 94.11 55.92 36.62 22.85 28.50 88.57 54.13 11.56
Qwen2.5-72B 95.86 62.53 35.55 33.85 35.75 94.69 41.05 13.95
Qwen3-32B 93.32 86.14 98.50 54.75 56.70 96.52 93.75 50.72
Qwen3-30B-A3B 93.14 85.20 98.15 43.89 58.25 75.36 92.85 44.11
Llama3.1-8B 88.58 9.05 62.15 3.30 8.80 37.46 45.15 2.74
Llama3.3-70B 94.74 49.26 94.72 40.30 41.75 96.53 88.95 37.25
Llama4-Scout-17B-16E 94.85 62.65 97.11 29.04 49.5 58.67 91.24 29.54
Mistral-Small-3.1-24B 94.65 28.96 49.55 24.55 36.65 66.98 77.55 16.87
Mistral-Large-123B 94.85 48.95 43.74 43.12 50.61 81.62 62.25 21.85
@@
GPT-4o 95.22 64.45 45.64 42.60 45.30 94.04 57.65 24.87
GPT-4.1 95.30 73.29 45.71 59.90 63.80 93.89 59.31 31.74
o4-mini 95.25 90.72 98.25 55.05 64.61 85.22 93.75 56.23
o3 95.02 89.12 98.45 71.51 78.62 90.97 91.44 64.87
DeepSeek-V3 95.12 65.25 98.35 68.65 71.10 96.55 93.43 45.74
DeepSeek-R1 94.92 93.61 98.15 62.50 69.80 89.54 97.92 57.14
Claude3.5-Sonnet 94.75 80.85 97.61 33.05 65.85 50.19 92.61 33.04
Claude3.7-Sonnet 94.85 74.18 98.11 57.43 66.35 86.51 93.53 57.91
Gemini2.5-Flash-Preview 94.69 89.72 98.12 35.85 50.35 71.20 93.33 38.69
Gemini2.5-Pro-Preview 94.86 91.45 97.95 54.80 76.90 71.26 94.15 56.25